MCGEE Company issued $650,000 face value bonds at a discount of $11,000. The bonds contain a call provision of 101. MCGEE decides to redeem the bonds due to a significant decline in interest rates. On that date, MCGEE had amortized only $2,500 of the discount.

Required:

1. Calculate the gain or loss on the early redemption of the bonds. Enter the amount as positive number. Round your answer to the nearest whole dollar.
$ Loss

2. Calculate the gain or loss on the redemption assuming that the call provision is 98 instead of 101. Enter the amount as positive number. Round your answer to the nearest whole dollar.
$ Gain

3. Select where the gain or loss should be presented on the financial statements.
Income Statement
Answer #1
1
Amount paid for redemption 656500 =650000*1.01
Less: Carrying value of bonds 641500 =650000-(11000-2500)
Loss on redemption of the bonds 15000
2
Amount paid for redemption 637000 =650000*0.98
Less: Carrying value of bonds 641500 =650000-(11000-2500)
Gain on redemption of the bonds 4500
3
On Income Statement